A. Write and balance the chemical equation for each of these reactions:


1. Solid calcium carbonate is synthesized from calcium oxide powder and carbon dioxide gas.

2. The reaction between zinc metal and hydrochloric acid forms solid zinc chloride and hydrogen gas.

3. The reaction between the aqueous solutions of silver nitrate and potassium chromate forms silver chromate precipitate and aqueous potassium nitrate.


1
Expert's answer
2021-03-18T02:58:28-0400

1. Ca(OH)2 + CO2 = CaCO3 + H2O.

2. Zn + 2HCl = ZnCl2 + H2.

3. K2Cr2O7 + 2AgNO3 → Ag2Cr2O7 + 2KNO3